1. Interchange the order of integration and evaluate the integral so Ix y dydx. 2. Use Lagrange multipliers to find the point question2
Answer #1

Answer: g= (4x²y , x) The Given curve is a is xy = 16 Let, g(x, y) = x²-16 g(x,y) = 0 f(x,y) = x² + y2 if = (2x, 24 Now, If =Putting the value oft in we e get 2x = 24 x3y * x = 4y? x 784% We get, Put the value of x2 in x y = 16 ya * (44²) y = 16 lby